OpenStudy (anonymous):

logarithmic equation? log2 (x+2) - log2 (x-5) =3

OpenStudy (saifoo.khan):

\[\log_2 (x+2) - \log_2 (x-5) = \log_2 2^3\]

OpenStudy (saifoo.khan):

Can you solve further?

OpenStudy (anonymous):

start with \[\log_2(\frac{x+2}{x-5})=3\] the rewrite in equivalent exponential form

OpenStudy (anonymous):

first equality because \[\log(A)-\log(B)=\log(\frac{A}{B})\]

OpenStudy (anonymous):

then go to \[\frac{x+2}{x-5}=2^3=8\] and then algebra from there on in